### v1

@@ 2.3 Effect of the ratio of labeled samples
Here, we compare the proposed FastECG and baseline models under various ratios of labeled samples in the datasets. Specifically, we adjust the ratio of the labeled samples in the dataset from 5% to 15% and present the averaged performance of different models on the four datasets in Fig. 3(a). The experiment results demonstrate the superiority of the proposed FastECG compared with FixMatch and FixMatch with LoRA under various ratios of the labeled data, especially when the ratio is low. As the ratio decreases from 15% to 5%, the performance advantage of FastECG over other models becomes more significant. When using 15% labeled data, FastECG achieves improvements of 1.3% on the macro $F_{\beta=2}$ compared to FixMatch with LoRA. In contrast, FastECG outperforms it by 1.9% on the macro $F_{\beta=2}$ using 5% labeled data. In Extend Data Fig.References, we also compare FastECG with other baseline models, where FastECG consistently outperforms them in CVDs detection under various labeled ratios.
Figure 3: The sensitivity analysis on the FastECG with the base backbone. (a) We adjust the ratio of the labeled samples in the dataset from 0.05 to 0.15 and report the averaged performance of different models across four datasets and six random seeds. (b) Averaged performance of the FastECG across four datasets and six random seeds under different initial ranks $r$. (c) We adjust the ratio of the important weight matrices to the total number of weight matrices and report the averaged performance across four datasets and six random seeds. Important weights are adapted with rank $r$ while the remaining weights are adapted with rank $\frac{1}{2}r$. (d) Averaged performance and training time of the FastECG across four datasets and six random seeds under different deactivation probability $p$.
@@ 2.4 Rank initialization in the one-shot rank allocation
Rank initialization is an important component in low-rank adaptation, which controls the number of trainable parameters during model training. In this section, we adjust the initial rank from 4 to 32 and present the averaged model performance on the four datasets in Fig.3(b). Note that the labeled ratio is set to 5%. The results indicate that FastECG with high initial ranks ($r=16,32$) achieves better performance than that with low initial ranks ($r=4,8$). This is because the model with higher ranks has more trainable parameters and thus demonstrates a larger capacity during training.
@@ 2.5 Effect of the number of important weight matrices
Based on the proposed one-shot rank allocation, FastECG allocates a rank $r$ to the incremental matrices with high importance and a rank $r/2$ to the matrices with low importance. The ratio of the important matrices to the total number of pre-trained matrices is defined as the coefficient $c$. The higher the coefficient is, the higher the ratio of the important matrices. In Fig.3(c), we adjust the coefficient from 0.2 to 0.8 and report the averaged model performance across four datasets. Note that the labeled ratio is set to 5%, and the initial ranks $r$ for all the low-rank matrices are set to 16. It can be observed that the performance of the proposed model is relatively insensitive to the changes in the $c$. In the Extended Data Fig.References, we visualize the rank distribution generated by the proposed method under various coefficients $c$. When the ratio of important matrices decreases from 0.8 to 0.2, the proposed method allocates more ranks to the self-attention and classification blocks than the convolution blocks. This phenomenon indicates that the deep modules exhibit higher importance than the shallow modules during model training, which aligns with the conclusions made by previous studies[42, 26].
@@ 2.6 Effect of the deactivation probability
For each pre-trained weight $W_{0}^{i}$ in the FastECG, the proposed RD-LoRA deactivates its low-rank matrices ($A^{i},B^{i}$) in the current iteration at a probability of $p$, which produces multiple sub-networks during model training. All the low-rank matrices are activated in the testing stage, generating an ensemble network that combines all the sub-networks. Consequently, the probability $p$ is an important parameter that controls the training time and the final performance of the proposed FastECG. In Fig.3(d), we adjust $p$ from 0.1 to 0.5 and present the averaged model performance across four datasets, including the training time for each iteration. Note that the labeled ratio is set to 5%, and the initial ranks for all the low-rank matrices are set to 16. The results show that the FastECG with $p=0.2$ demonstrates the best detection performance compared with the model with other settings. In addition, it can be observed that the training time of the FastECG decreases as $p$ increases. The reason is that the larger the $p$ is, the more low-rank matrices are deactivated during model training, which speeds up the forward-backward propagation.
2.7 Performance comparisons under different backbone sizes
@@
Previous studies have demonstrated that the backbones pre-trained on a large-scale dataset with a large model size usually have better performance than the other backbones[43, 44]. In the previous sections, we have already proved the robustness and computation efficiency of the proposed FastECG under a base backbone with 9.505 million parameters. Here, we compare its performance with other baseline models under medium and large backbones, which share the same architecture as the base backbone but have more parameters. Specifically, the medium backbone has 50.494 million parameters, and the large backbone has 113.490 million parameters (Extended Data Table Computation-Efficient Semi-Supervised Learning for ECG-based Cardiovascular Diseases Detection). They are pre-trained on the CODE-full dataset, a huge but restricted ECG dataset with 2,322,513 ECG recordings from 1,558,772 patients[31, 5]. In Extended Data Table Computation-Efficient Semi-Supervised Learning for ECG-based Cardiovascular Diseases Detection and Table Computation-Efficient Semi-Supervised Learning for ECG-based Cardiovascular Diseases Detection, we report the performance of FastECG and semi-supervised baselines on the medium and the large backbones, respectively. The results demonstrate that FastECG achieves similar and even better CVDs detection performance than the semi-supervised baselines and exhibits the lowest computation costs. For example, using the medium backbone, FastECG achieves a macro $F_{\beta=2}$ of 0.599$\pm$0.010, which is 3.7% larger than the second-best model’s (SAW) performance in the PTB-XL dataset. Using the large backbone, FastECG achieves a macro $F_{\beta=2}$ of 0.565$\pm$0.010 in the G12EC dataset, outperforming SAW by 3.1%. Regarding the computational costs, the number of trainable parameters of FastECG is 0.9% to 3.1% of the other baselines on the medium backbone and 0.6% to 2.1% on the large backbone. In addition, FastECG demonstrates the lowest GPU memory consumption and the highest training speed compared to the other semi-supervised baselines. For the memory footprint, FastECG achieves an average GPU memory usage of 6.16 GB using the medium backbone and 9.22 GB using the large backbone, 3.09 GB and 4.59 GB less than the second-best model (Adsh). Furthermore, FastECG achieves an average training time per iteration of 259.25 ms using the medium backbone and 485.5 ms using the large backbone, 162.5 ms and 289.75 ms faster than the second-best model (MixedTeacher). These phenomenons demonstrate that as the number of model parameters increases, the computational efficiency advantage of FastECG over other models becomes increasingly apparent. In Extended Data Table Computation-Efficient Semi-Supervised Learning for ECG-based Cardiovascular Diseases Detection and Table Computation-Efficient Semi-Supervised Learning for ECG-based Cardiovascular Diseases Detection, we present the performance of FastECG and parameter-efficient semi-supervised methods on the medium and large backbones, respectively. It can be observed that FastECG outperforms the other models in CVDs detection on both medium and large backbones. Additionally, FastECG demonstrates the fastest training speed across four datasets compared with other parameter-efficient methods. In Supplementary Materials Section 3 Fig. S2-S3, we provide the paired t-test results of the model performance on the two backbones. The statistical results indicate that FastECG outperforms the above baselines in ECG-based CVDs detection at a 0.05 significance level in most conditions.
@@ 2.8 Ablation study
We conducted an ablation study to evaluate the contribution of the modules implemented in FastECG. Specifically, we remove each of the proposed three modules from the FastECG and record the corresponding model performance on the four datasets. Note that the initial rank $r$ is 16 for all the compared models. (1) The random-deactivation low-rank adaptation improves model performance and computational efficiency. It randomly deactivates the low-rank matrices in each trainable layer with a given probability, which generates multiple sub-networks during model training. All the low-rank matrices are activated in the testing stage, and the sub-networks are merged into a robust ensemble network. In the G12EC dataset, the FastECG with the random deactivation module increases the macro $F_{\beta=2}$ from 0.536$\pm$0.021 to 0.551$\pm$0.017 and decreases the ranking loss from 0.095$\pm$0.004 to 0.092$\pm$0.002. Additionally, the Time/iter is reduced from 104 ms to 98 ms. This observation indicates that the computation cost of training a sub-network with fewer low-rank matrices is lower than optimizing the entire network. (2) The one-shot rank allocation method improves low-rank adaptation performance with high computational efficiency. One drawback of LoRA[25] is its inability to allocate an optimal rank for each incremental matrix[26, 30]. Here, the proposed rank allocation method determines the optimal ranks using only one gradient-backward iteration with high computational efficiency. As shown in Table 2, removing the proposed module decreases the detection performance of FastECG in the four datasets. For example, the macro $F_{\beta=2}$ decreases from 0.530$\pm$0.008 to 0.514$\pm$0.018 and the coverage increases from 2.483$\pm$0.055 to 2.503$\pm$0.040 on the Chapman dataset. Additionally, it can be observed that the one-shot rank allocation module does not introduce heavy computational burdens (Time/iter only increases by 1-2ms), demonstrating its high computational efficiency. (3) The proposed lightweight semi-supervised learning benefits model performance under limited supervision without greatly increasing the training time. It utilizes the unlabeled data to stabilize the statistics within the BN layers in the convolution blocks, preventing them from over-fitting to small amounts of labeled data. When the module is removed, the performance of the FastECG decreases on all the datasets. For example, the macro $F_{\beta=2}$ score decreases from 0.551$\pm$0.017 to 0.536$\pm$0.029 on the G12EC dataset. Compared to other semi-supervised baselines (Table 1), the extra computation costs (Time/iter only increases by 20ms) caused by the proposed module are much lower. In Supplementary Materials Section 3, ablation studies on the medium and large backbones are provided, consistently demonstrating the notable contribution of the proposed three techniques.
@@
Table 2: Ablation study of the proposed FastECG on the base backbone. ’w/o random deactivation’ represents the FastECG without the random deactivation technique, and the deactivation probability $p$ is set to zero. ’w/o rank allocation’ represents the FastECG without the one-shot rank allocation, and all pre-trained weights are updated with the initial rank $r=16$. ’w/o unlabeled data’ denotes the FastECG without the unlabeled data for lightweight semi-supervised learning.
@@
w/o unlabeled data 77ms 0.040$\pm$0.002 2.468$\pm$0.050 0.896$\pm$0.010 0.533$\pm$0.010 0.350$\pm$0.020 0.527$\pm$0.026
FastECG 97ms 0.040$\pm$0.002 2.483$\pm$0.055 0.896$\pm$0.006 0.536$\pm$0.004 0.355$\pm$0.005 0.530$\pm$0.008
3 Discussion
@@
Bottlenecks in model performance and computational efficiency have become great challenges in the clinical application of CVDs detection systems based on pre-trained models, especially when the supervised information is scarce in the downstream ECG datasets. Previous studies usually overcome the performance bottleneck at the cost of a large drop in computational efficiency[18, 26, 30, 22]. In this paper, we propose a computationally efficient semi-supervised learning paradigm (FastECG) for adapting the PMs on downstream datasets with limited supervision and high computational efficiency. Experiment results on four downstream ECG datasets and three backbone settings indicate that FastECG achieves superior CVDs detection performance and computational efficiency compared to state-of-the-art methods.
Our proposed FastECG paradigm enhances the performance of pre-trained models without a large sacrifice of computational efficiency, including GPU memory footprint, training time, and storage consumption. Through integrating the sub-networks generated by the random-deactivation operation, the expected CVDs detection performance of FastECG is superior to the other compared methods, such as LoRA[25] and DyLoRA[40]. Its training time is also decreased due to omitting the computation of the update matrices for the deactivated low-rank matrices. In addition, determining the optimal rank distribution is a popular method to improve the low-rank adaptation performance of the pre-trained models. However, AdaLoRA[26] and IncreLoRA[30] continuously adjust the optimal ranks during the training process, decreasing the model computation efficiency. In contrast, the one-shot allocation method avoids this by determining the optimal ranks using only one forward-backward propagation before model training. In the downstream ECG datasets, the label scarcity problem is a crucial factor that results in the performance bottleneck of pre-trained models. Although semi-supervised learning offers an effective solution to address the label scarcity problem and overcome the bottleneck[45, 21, 46, 24], the commonly utilized semi-supervised techniques such as consistency training can be expensive in large pre-trained models and greatly increase their training time[47]. In contrast, FastECG adopts a lightweight semi-supervised learning pipeline, which leverages the unlabeled data to stabilize the statistic estimation in batch normalization layers. The extracted features of unlabeled data can be released in the self-attention blocks, which saves a lot of GPU memory footprint and computation time. Our results demonstrate that the lightweight semi-supervised learning method can help FastECG overcome the performance bottleneck without significantly decreasing its computational efficiency. In summary, the theoretical and empirical analysis indicates that FastECG is able to improve the detection performance and computational efficiency simultaneously.
In conclusion, our study offers a fast and robust semi-supervised learning paradigm for ECG-based CVDs detection under limited supervision. It provides a feasible solution for efficiently adapting pre-trained models on downstream ECG datasets. We hope this learning paradigm will pave the way for the application of automatic CVDs detection systems and broaden its applicability to various ECG-based tasks.

### v2

@@ 5.2 Effect of the Ratio of Labeled Samples
Here, we compare the proposed CE-SSL and baseline models under various ratios of labeled samples in the datasets. Specifically, we adjust the ratio of the labeled samples in the dataset from 5% to 15% and present the averaged performance of different models on the four datasets in Fig. 3. The experiment results demonstrate the superiority of the proposed CE-SSL compared with FixMatch and FixMatch with LoRA under various ratios of the labeled data, especially when the ratio is low. As the ratio decreases from 15% to 5%, the performance advantage of CE-SSL over other models becomes more significant. When using 15% labeled data, CE-SSL achieves improvements of 1.3% on the macro $F_{\beta=2}$ compared to FixMatch with LoRA. In contrast, CE-SSL outperforms it by 1.9% on the macro $F_{\beta=2}$ using 5% labeled data. In Supplementary Materials Fig. S4, we also compare CE-SSL with other baseline models, where CE-SSL consistently outperforms them in CVDs detection under various labeled ratios.
Figure 3: Effect of the ratio of labeled samples for model training. We adjust the ratio of the labeled samples in the dataset from 0.05 to 0.15 and report the averaged performance of different models across four datasets and six random seeds.
@@ 5.3 Rank Initialization in the One-Shot Rank Allocation
Rank initialization is an important component in low-rank adaptation, which controls the number of trainable parameters during model training. In this section, we adjust the initial rank from 4 to 32 and present the averaged model performance on the four datasets in Fig.4. Note that the labeled ratio is set to 5%. The results indicate that CE-SSL with high initial ranks ($r=16,32$) achieves better performance than that with low initial ranks ($r=4,8$). This is because the model with higher ranks has more trainable parameters and thus demonstrates a larger capacity during training.
Figure 4: Effect of the rank initialization. Averaged performance of the CE-SSL across four datasets and six random seeds under different initial ranks $r$.
@@ 5.4 Effect of the Number of Important Weight Matrices
Based on the proposed one-shot rank allocation, CE-SSL allocates a rank $r$ to the incremental matrices with high importance and a rank $r/2$ to the matrices with low importance. The ratio of the important matrices to the total number of pre-trained matrices is defined as the coefficient $c$. The higher the coefficient is, the higher the ratio of the important matrices. In Fig.5, we adjust the coefficient from 0.2 to 0.8 and report the averaged model performance across four datasets. Note that the labeled ratio is set to 5%, and the initial ranks $r$ for all the low-rank matrices are set to 16. It can be observed that the performance of the proposed model is relatively insensitive to the changes in the $c$. In Supplementary Materials Fig. S5, we visualize the rank distribution generated by the proposed method under various coefficients $c$. When the ratio of important matrices decreases from 0.8 to 0.2, the proposed method allocates more ranks to the self-attention and classification blocks than the convolution blocks. This phenomenon indicates that the deep modules exhibit higher importance than the shallow modules during model training, which aligns with the conclusions made by previous studies[53, 28].
Figure 5: Effect of the ratio of important weight matrices. We adjust the ratio of the important weight matrices to the total number of weight matrices and report the averaged performance across four datasets and six random seeds. Important weights are adapted with rank $r$ while the remaining weights are adapted with rank $\frac{1}{2}r$.
@@ 5.5 Effect of the Deactivation Probability
For each pre-trained weight $W_{0}^{i}$ in the CE-SSL, the proposed RD-LoRA deactivates its low-rank matrices ($A^{i},B^{i}$) in the current iteration at a probability of $p$, which produces multiple sub-networks during model training. All the low-rank matrices are activated in the testing stage, generating an ensemble network that combines all the sub-networks. Consequently, the probability $p$ is an important parameter that controls the training time and the final performance of the proposed CE-SSL. In Fig.6, we adjust $p$ from 0.1 to 0.5 and present the averaged model performance across four datasets, including the training time for each iteration. Note that the labeled ratio is set to 5%, and the initial ranks for all the low-rank matrices are set to 16. The results show that the CE-SSL with $p=0.2$ demonstrates the best detection performance compared with the model with other settings. In addition, it can be observed that the training time of the CE-SSL decreases as $p$ increases. The reason is that the larger the $p$ is, the more low-rank matrices are deactivated during model training, which speeds up the forward-backward propagation.
Figure 6: Effect of the deactivation probability.Averaged performance and training time of the CE-SSL across four datasets and six random seeds under different deactivation probability $p$.
5.6 Performance Comparisons under Various Backbone Sizes
@@
In the previous sections, we have already proved the robustness and computation efficiency of the proposed CE-SSL under a base backbone with 9.505 million parameters. Here, we compare its performance with other baseline models under medium and large backbones, which share the same architecture as the base backbone but have more parameters (Table I). Specifically, the medium backbone has 50.494 million parameters, and the large backbone has 113.490 million parameters. They are pre-trained on the CODE-full dataset, a huge but restricted ECG dataset with 2,322,513 ECG recordings from 1,558,772 patients[39, 5]. In Supplementary Materials Table S9 and Table S10, we report the performance of CE-SSL and semi-supervised baselines on the medium and the large backbones, respectively. The results demonstrate that CE-SSL achieves similar and even better CVDs detection performance than the semi-supervised baselines and exhibits the lowest computation costs. For example, using the medium backbone, CE-SSL achieves a macro $F_{\beta=2}$ of 0.599$\pm$0.010, which is 3.7% larger than the second-best model’s (SAW) performance in the PTB-XL dataset. Using the large backbone, CE-SSL achieves a macro $F_{\beta=2}$ of 0.565$\pm$0.010 in the G12EC dataset, outperforming SAW by 3.1%. Regarding the computational costs, the number of trainable parameters of CE-SSL is 0.9% to 3.1% of the other baselines on the medium backbone and 0.6% to 2.1% on the large backbone. In addition, CE-SSL demonstrates the lowest GPU memory consumption and the highest training speed compared to the other semi-supervised baselines. For the memory footprint, CE-SSL achieves an average GPU memory usage of 6.16 GB using the medium backbone and 9.22 GB using the large backbone, 3.09 GB and 4.59 GB less than the second-best model (Adsh). Furthermore, CE-SSL achieves an average training time per iteration of 259.25 ms using the medium backbone and 485.5 ms using the large backbone, 162.5 ms and 289.75 ms faster than the second-best model (MixedTeacher). These phenomenons demonstrate that as the number of model parameters increases, the computational efficiency advantage of CE-SSL over other models becomes increasingly apparent. In Supplementary Materials Table S11 and Table S12, we present the performance of CE-SSL and parameter-efficient semi-supervised methods on the medium and large backbones, respectively. It can be observed that CE-SSL outperforms the other models in CVDs detection on both medium and large backbones. Additionally, CE-SSL demonstrates the fastest training speed across four datasets compared with other parameter-efficient methods. In Supplementary Materials Section 3 Fig. S2-S3, we provide the paired t-test results of the model performance on the two backbones. The statistical results indicate that CE-SSL outperforms the above baselines in ECG-based CVDs detection at a 0.05 significance level in most conditions.
5.7 Toward Higher Computational Efficiency in Clinical Practices
Although deploying the CE-SSL paradigm with the base backbone on low-level devices (4-6 GB GPU memory) is easy, implementing the paradigm with the medium and large backbones is still challenging. To overcome this limitation, we adopt a simple but effective approach to boost the computational efficiency of the CE-SSL. Specifically, we freeze the first two convolution blocks in the backbones during the CE-SSL training process. The new paradigm is denoted as ’CE-SSL-F’ in the following analysis. We present the CVDs detection performance and the computational efficiency of CE-SSL-F, CE-SSL, and the SOTA methods in semi-supervised learning in Fig.7. Note that the batch sizes for all the compared methods are set to 64.
First, freezing the convolution blocks greatly reduces the cached activation during the forward pass, significantly decreasing the GPU memory footprints. As shown in Fig.7a, it can be observed that the CE-SSL-F requires nearly 50% less GPU memory footprints compared to the CE-SSL, generalizing its applications in low-level devices (NVIDIA RTX 3050 laptops and RTX 4060 GPU cards). Specifically, CE-SSL-F is deployable on RTX 3050 laptops with both base and medium backbones, and it is the only method that can be implemented on the RTX 4060 GPU cards with a large backbone. In contrast, deploying the CE-SSL with a large backbone requires medium-level devices (NVIDIA RTX 4070 GPU cards), while other semi-supervised methods require high-level devices with GPU memory larger than 12 GB. Second, the parameters of the frozen blocks are not updated during the backward pass, which increases the training speed of CE-SSL-F. The larger the backbone is, the more parameters are frozen, and thus the more gradient backward time is saved. As shown in Fig.7b, CE-SSL-F demonstrates the fastest training speed compared with other models, and its advantages become more significant along with the increase in backbone sizes. Third, CE-SSL-F only sacrifices 1-2% CVDs detection performance compared with CE-SSL. More importantly, it consistently outperforms the other semi-supervised methods across different backbones (Fig.7c), demonstrating its effectiveness in CVDs detection. This phenomenon can be explained by the strong transferability of the pre-trained convolution blocks located in the first few layers of the backbone[54, 55]. Specifically, they mainly contain domain-invariant knowledge for CVDs detection, and their parameters will not be changed significantly during the fine-tuning process. Therefore, freezing them does not greatly decrease the model performance.
In summary, the experiment results illustrate that the computational efficiency of the CE-SSL can be increased to adapt to low-level devices without losing its superior CVDs detection performance compared to other semi-supervised methods. This advantage demonstrates CE-SSL’s flexibility in different clinical application scenarios with various computational resources.
(a) The peak GPU memory footprints during training
(b) The average training time for each optimization iteration
(c) Macro $F_{\beta=2}$ score
Figure 7: Evaluation of the average detection performance and computational efficiency of various semi-supervised methods for CVDs across four downstream datasets. Specifically, the macro $F_{\beta=2}$ scores are used to evaluate their CVDs detection performance. Additionally, the average training time per optimization iteration and the maximum GPU memory usage are presented to evaluate computational efficiency. Methods exceeding the GPU memory thresholds, indicated by the red dashed lines, are not deployable on the corresponding NVIDIA GPU cards. The only distinction between CE-SSL-F and CE-SSL lies in the training process. Specifically, CE-SSL-F freezes the first two convolution blocks during training, whereas CE-SSL does not impose such constraints on them. The initial rank for CE-SSL and CE-SSL-F is set to 16 and 4, respectively.
@@ 5.8 Ablation Study
We conducted an ablation study to evaluate the contribution of the modules implemented in CE-SSL. Specifically, we remove each of the proposed three modules from the CE-SSL and record the corresponding model performance on the four datasets. Note that the initial rank $r$ is 16 for all the compared models. (1) The random-deactivation low-rank adaptation improves model performance and computational efficiency. It randomly deactivates the low-rank matrices in each trainable layer with a given probability, which generates multiple sub-networks during model training. All the low-rank matrices are activated in the testing stage, and the sub-networks are merged into a robust ensemble network. In the G12EC dataset, the CE-SSL with the random deactivation module increases the macro $F_{\beta=2}$ from 0.536$\pm$0.021 to 0.551$\pm$0.017 and decreases the ranking loss from 0.095$\pm$0.004 to 0.092$\pm$0.002. Additionally, the Time/iter is reduced from 104 ms to 98 ms. This observation indicates that the computation cost of training a sub-network with fewer low-rank matrices is lower than optimizing the entire network. (2) The one-shot rank allocation method improves low-rank adaptation performance with high computational efficiency. One drawback of LoRA[27] is its inability to allocate an optimal rank for each incremental matrix[28, 32]. Here, the proposed rank allocation method determines the optimal ranks using only one gradient-backward iteration with high computational efficiency. As shown in Table III, removing the proposed module decreases the detection performance of CE-SSL in the four datasets. For example, the macro $F_{\beta=2}$ decreases from 0.530$\pm$0.008 to 0.514$\pm$0.018, and the coverage increases from 2.483$\pm$0.055 to 2.503$\pm$0.040 on the Chapman dataset. Additionally, it can be observed that the one-shot rank allocation module does not introduce heavy computational burdens (Time/iter only increases by 1-2ms), demonstrating its high computational efficiency. (3) The proposed lightweight semi-supervised learning benefits model performance under limited supervision without greatly increasing the training time. It utilizes the unlabeled data to stabilize the statistics within the BN layers in the convolution blocks, preventing them from over-fitting to small amounts of labeled data. When the module is removed, the performance of the CE-SSL decreases on all the datasets. For example, the macro $F_{\beta=2}$ score decreases from 0.551$\pm$0.017 to 0.536$\pm$0.029 on the G12EC dataset. Compared to other semi-supervised baselines (Table II), the extra computation costs (Time/iter only increases by 20ms) caused by the proposed module are much lower. In Supplementary Materials Section 3, ablation studies on the medium and large backbones are provided, consistently demonstrating the notable contribution of the proposed three techniques.
@@
TABLE III: Ablation study of the proposed CE-SSL on the base backbone. ’w/o random deactivation’ represents the CE-SSL without the random deactivation technique, and the deactivation probability $p$ is set to zero. ’w/o rank allocation’ represents the CE-SSL without the one-shot rank allocation, and all pre-trained weights are updated with the initial rank $r=16$. ’w/o semi-supervised BN’ denotes the CE-SSL without the semi-supervised batch normalization for lightweight semi-supervised learning.
@@
w/o semi-supervised BN 77ms 0.040$\pm$0.002 2.468$\pm$0.050 0.896$\pm$0.010 0.533$\pm$0.010 0.350$\pm$0.020 0.527$\pm$0.026
CE-SSL 97ms 0.040$\pm$0.002 2.483$\pm$0.055 0.896$\pm$0.006 0.536$\pm$0.004 0.355$\pm$0.005 0.530$\pm$0.008
6 Conclusion
@@
Bottlenecks in model performance and computational efficiency have become great challenges in the clinical application of CVDs detection systems based on pre-trained models, especially when the supervised information is scarce in the downstream ECG datasets. Previous studies usually overcome the performance bottleneck at the cost of a large drop in computational efficiency[15, 28, 32, 19]. In this paper, we propose a computationally efficient semi-supervised learning paradigm (CE-SSL) for adapting pre-trained models on downstream datasets with limited supervision and high computational efficiency. Experiment results on four downstream ECG datasets and three backbone settings indicate that CE-SSL achieves superior CVDs detection performance and computational efficiency compared to state-of-the-art methods. In conclusion, our study offers a fast and robust semi-supervised learning paradigm for ECG-based CVDs detection under limited supervision. It provides a feasible solution for efficiently adapting pre-trained models on downstream ECG datasets. We hope this learning paradigm will pave the way for the application of automatic CVDs detection systems and broaden its applicability to various ECG-based tasks.
7 Limitation and Future Work